Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
berni - 15:56 Tuesday 07 March 2023 (59161) Print this report
Operator Report - Morning shift

ITF foun in CARM_NULL_1F; for this reason I skipped the OSD calibration.

At 7:00 started the planned Maintenance, below the list of activities comunicated to the control room:

  • standard vacuum and cleaning operations;
  • UTA DET cold coil removal;
  • ENV-TCSR2 Slowmon box added;
  • PSL: neovan recovery (still in progress);
  • DAQ: demodulation server maintenance;
  • at 10:42 Valerio went at WE to perform some test on the LC;
  • work by electronics team in TCS room to check the problem at the flip mirrors;
  • test on PduServer by Franco and Gkaitatzis

 

DET

At around 7:00 UTC I started the procedure to lock and scan the OMC; unfurtunately it was interrupted because of the work at the PSL.

 

SUSP

MC LC open by the guardian; properly closed at 14:20 UTC. The alignment must be checkes once the beam will be available again.

 

HVAC

At 7:09 UTC I put the UTA INJ in "Portata Nominale"

 

DAQ-Computing

The DMS is showing a red flag for DISK - DAQ_DISK_FULL..DATADEVWEB reporting that data/dev/web is above 95 percent full

 

Pending actions

DET
( - ) Restore the floating setpoint for SDB1_LC

Images attached to this report
Comments to this report:
bersanetti - 18:04 Tuesday 07 March 2023 (59164) Print this report

During the maintenance the commands which enable the Central Heating have been commented (lines 381 and 383 of ITF_LOCK.py), in order not to spoil the beginning of the TCS activity. This will need to be reverted before the ITF recovery.

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×